A zero-order reaction has a constant rate of 4.90×10−4M/s. If after 50.0 seconds the concentration has dropped to 8.50×10−2M, what was the initial concentration?
Answer #1

if zero roder

RATE = K

RATE = 4.9*10^-4

t = 50 s

then

A = A0 - kt

(8.5*10^-2) = A0 -( 4.9*10^-4) * 50

A0 = (8.5*10^-2) +( 4.9*10^-4) * 50

A0 = 0.1095 M